# Lattice-induced sound trapping in biperiodic metasurfaces of acoustic resonators

### v3

@@ II.3 Selection rules for multipole coupling in 2D lattices
The further consideration is facilitated by the symmetry analysis and associated selection rules that constrain which multipoles interact in structures with a given symmetry [52, 63, 47], or in other words, which off-diagonal elements of $\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ in Eq. (10) do not vanish.
@@
Let us first consider the acoustic interaction between scalar multipoles induced by a 2D Bravais lattice with rotational symmetry of order $\mathfrak{n}$ in the $xy$ plane. We also assume that each unit cell contains only one sphere, for which the single-particle T-matrix is diagonal $T_{\ell,m,\ell^{\prime},m^{\prime}}=\mathfrak{a}_{\ell}\delta_{\ell,\ell^{\prime}}\delta_{m,m^{\prime}}$ (see Appendix C for the definition of $\mathfrak{a}_{\ell}$). Hence, the diagonal elements of $\mathbf{T}^{-1}_{\rm eff}$ become $\left(1/T_{\ell,m,\ell,m}-\Sigma^{\prime}_{\ell,m,\ell,m}\right)$, while the off-diagonal elements are $\left(-\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}\right)$ with $\ell\neq\ell^{\prime}$ and $m\neq m^{\prime}$. Thus, the unit cell has inversion symmetry (considering $\mathbf{r}\to-\mathbf{r}$) and rotational symmetry of order $\mathfrak{n}$ (considering rotations by $2\pi/\mathfrak{n}$ around the $z$ axis). Lattice-induced coupling between two different multipoles $(\ell,m)$ and $(\ell^{\prime},m^{\prime})$ is allowed ($\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}\neq 0$) if the multipoles have the same symmetry with respect to the symmetry operations upon which the system repeats itself. For an arbitrary $\mathbf{k}_{\parallel}$, the selection rule is $\left(\ell+\ell^{\prime}\right)=(m-m^{\prime})\ \text{mod}\ 2$ [64]. Under normal incidence $\mathbf{k}_{\parallel}=\bm{0}$, the following conditions must also be met (see Appendix E)
$\ell=\ell^{\prime}\ \text{mod}\ 2\,,\quad\text{and}\quad m=m^{\prime}\ \text{mod}\ \mathfrak{n}\,.$ (12)
@@
Otherwise, the lattice-induced coupling is zero, and off-diagonal entries $\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}$ in Eq. (9) vanish. The first condition states that the multipoles must have the same parity, which follows from the inversion symmetry of the lattice. In contrast, the second condition follows from the rotational symmetry of the lattice unit cell and the incident wave.
However, even though $\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}=0$, the particle-induced coupling is still present in non-spherical acoustic resonators due to off-diagonal entries of the T-matrix $T_{\ell,m,\ell^{\prime},m^{\prime}}\neq 0$, which can be nonzero only if the indices obey the same rules (12) (in this case, $\mathfrak{n}$ will be the order of rotational symmetry of an isolated resonator) [52]. In the general case, $\mathfrak{n}$ should be chosen as the minimum of rotational symmetry orders for the unit cell and the resonator.
@@ II.4 Reflection and transmission coefficients
Effective multipole moments (11) allow us to calculate the reflection and transmission coefficients of a lattice for an acoustic plane wave. To do this, we consider the pressure field in the far-field zone for $z<0$ and $z>0$. The incident wave is a pressure plane wave $p_{\rm inc}(\omega;\mathbf{r})=p_{0}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}$ with $\mathbf{k}=\mathbf{k}_{\parallel}+\sqrt{k^{2}-k_{\parallel}^{2}}\hat{\mathbf{z}}$ and $|\mathbf{k}|=k$. Using Eqs. (3) and (5) with multipole coefficients (11), one can write the total pressure after the summation over the lattice sites as
@@
Note that the second term in the brackets in (13) corresponds to the scattered field in the presence of an external field $p_{0}\neq 0$ and the field of a lattice eigenmode in the absence $p_{0}=0$ (see Sec. III.1).
If the wavelength in the background medium with $c_{\rm b}$ is larger than the lattice constant $L$, lattice field sums (14) can be written in the far-field approximation as (see Appendix F)
$\Sigma_{\ell,m}(\omega,\mathbf{k}_{\parallel};\mathbf{r})\approx\frac{2\pi}{k^{2}A}(-\mathrm{i})^{\ell}Y_{\ell,m}(\theta_{\mathbf{k}_{\pm}},\varphi_{\mathbf{k}_{\pm}})\mathrm{e}^{\mathrm{i}\mathbf{k}_{\pm}\cdot\mathbf{r}}\,$, (15)
where the wave vectors are $\mathbf{k}_{\pm}=\mathbf{k}_{\parallel}\pm\sqrt{k^{2}-k_{\parallel}^{2}}\hat{\mathbf{z}}$ for $z\gtrless 0$ and $A$ is the area of a lattice unit cell ($A=L^{2}$ for a square unit cell). By inserting (15) into Eq. (13), we obtain the pressure field reflection and transmission coefficients for a lattice as
@@
The quantities in (16a)-(17) depend on $\omega$ and $\mathbf{k}_{\parallel}$ due to $a_{\ell,m}^{\rm eff}\equiv a_{\ell,m}^{\rm eff}(\omega,\mathbf{k}_{\parallel})$. In (17), $\mathbf{k}_{+}$ points into the direction of propagation of the incident and transmitted plane waves, while $\mathbf{k}_{-}$ is the wave vector of the reflected wave. Thus, a subwavelength 2D lattice can radiate only in two directions $\mathbf{k}_{-}$ and $\mathbf{k}_{+}$ with the amplitude and phase dictated by $\mathcal{P}_{-}$ and $\mathcal{P}_{+}$, respectively. The response of the lattice is determined by the effective multipole moments, i.e., the renormalized response of one unit cell (11). Moreover, the lattice coupling does not change the angular dependence of the multipoles of a unit cell, given by $Y_{\ell,m}(\theta_{\mathbf{k}_{\pm}},\varphi_{\mathbf{k}_{\pm}})$, but only changes the multipole moments from (7) to (11). Thus, one can control the radiation pattern of the unit cell by tailoring the values of the multipole coefficients. Note also that, in the absence of the incident field $\mathbf{b}=\bm{0}$, expressions (17) describe the far-field amplitudes of the scattered waves radiated by a lattice eigenmode (see Sec. III.1).
Finally, let us consider the case of normal incidence $\mathbf{k}_{\parallel}=\bm{0}$. For the arguments $\cos\theta_{\pm}=\pm 1$ corresponding to $\theta=0$ (+z+z direction) and $\theta=\pi$ (z-z direction), the associated Legendre polynomials (30) take the following values [31]
$P^{m}_{\ell}(\pm 1)$ $=\begin{cases}0,&m\neq 0\\ (\pm 1)^{\ell},&m=0\end{cases}\,$, (18)
From Eq. (17), we see that for $\mathbf{k}_{\parallel}=\bm{0}$, the amplitudes of the scattered (radiated) fields separate into two subseries corresponding to even and odd multipoles, in full agreement with the first selection rule from (12),
@@
where $\mathcal{P}_{\pm}\equiv\mathcal{P}_{\pm}(\omega,\bm{0})$ and $a_{\ell,0}^{\rm eff}\equiv a_{\ell,0}^{\rm eff}(\omega,\bm{0})$ for compactness. Thus, the response of a lattice at normal incidence can be considered as a superposition of the responses of two noninteracting sublattices, each of which includes either only even multipoles or only odd multipoles, respectively.
III BIC in 2D lattices of spherical resonators
@@ III.1 Eigenmode and BIC condition
According to Cramer’s theorem, nontrivial solutions to homogeneous equation (20) exist if the inverse effective T-matrix is singular, i.e., $\det\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})=0$ [33].
For an open system and real $\mathbf{k}_{\parallel}$, a nontrivial solution to Eq. (20) can generally be found for a complex frequency $\omega(\mathbf{k}_{\parallel})-\mathrm{i}\gamma_{\rm r}(\mathbf{k}_{\parallel})-\mathrm{i}\gamma_{\rm nr}(\mathbf{k}_{\parallel})]$, where $\gamma_{\rm r}(\mathbf{k}_{\parallel})$ and $\gamma_{\rm nr}(\mathbf{k}_{\parallel})$ define energy dissipation due to radiative and nonradiative losses, respectively. However, photonic and acoustic 2D lattices can support eigenmodes without radiative losses $(\gamma_{\rm r}=0)$, which are just called bound states in the continuum or BICs. Moreover, in the absence of nonradiative losses (absorption), BICs emerge for pure real-valued frequencies due to $\gamma_{\rm nr}=0$. Hence, a criterion of the BIC occurrence can be formulated as
$\det\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})=0\,,\quad\text{and}\quad\omega\in\mathbb{R}\,.$ (21)
Since the matrix $\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ can be subject to the singular value decomposition (SVD), the absolute value of its determinant equals to the product of its singular values, which are nonnegative real-valued numbers [65]. Thus, if a singular value of the matrix is zero, then the determinant is also zero. Hence, the solution to Eq. (20) is the right singular vector of matrix $\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ with a singular value equal to zero [61, 66]. After denoting the operator that picks the smallest singular value of a matrix with $s_{\rm min}$, BIC condition (21) reads as
$s_{\rm min}\left[\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})\right]=0\,,\quad\text{and}\quad\omega\in\mathbb{R}\,.$ (22)
Conditions (21) and (22) ensure the existence of a lattice eigenmode with a real eigenfrequency, i.e., a BIC. If this condition is satisfied, the right singular vector $\mathbf{a}^{\rm eff}(\omega,\mathbf{k}_{\parallel})$ corresponding to the eigenfrequency $\omega=\omega_{\rm BIC}$ contains the multipole coefficients $a_{\ell,m}^{\rm eff}(\omega,\mathbf{k}_{\parallel})$. Inserting them into Eqs. (17) or (19) must give us that the far-field radiation amplitudes for the BIC are $\mathcal{P}_{+}=\mathcal{P}_{-}=0$.
@@ III.2 Even BIC in the monopole-quadrupole approximation
Equation (22) can provide us with resonance frequencies of BICs in arbitrary systems. Here, we give an example of the matrix $\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ for which Eq. (22) has a solution with $\mathbf{k}_{\parallel}=\bm{0}$ and a real eigenfrequency corresponding to the acoustic $\Gamma$-point BIC. For this mode, sound radiation in the $\pm z$ directions is suppressed, so that the lattice does not radiate at all, and the mode can be referred to as the accidental $\Gamma$-point BIC. According to selection rules (12), eigenmodes (including BICs) of a lattice of spheres can have either even or odd parity, which also determines their multipolar content. In the following, we choose the case of even parity. However, the odd-parity modes, which can also enable the BIC, can be considered in the same manner (see Appendix I).
Let us consider the acoustic eigenmode in the square lattice $\mathfrak{n}=4$ in (12)] with even parity, for which the following multipole coefficients can be nonzero
@@
where $2\ell^{\prime}=0,2,4,$... and $4m^{\prime}=0,\pm 4,...,\pm 2\ell^{\prime}$. The other multipole contributions must be zero according to selection rules (12). Furthermore, we also consider spherical scatterers (resonators) whose acoustic response in the given frequency range can be sufficiently described by the zonal monopole (M), dipole (D), quadrupole (Q), and octupole (O). At the same time, higher-degree multipoles can be neglected (see Fig. 2).
@@
In this case, we can keep only two terms, the zonal monopole $a_{\rm M}^{\rm eff}\equiv a_{0,0}^{\rm eff}$ and quadrupole $a_{\rm Q}^{\rm eff}\equiv a_{2,0}^{\rm eff}$, because only these multipoles have even parity. For $\mathbf{k}_{\parallel}=\bm{0}$, all unit cells in the array will have identical zonal monopole $a_{\rm M}^{\rm eff}$ and quadrupole $a_{\rm Q}^{\rm eff}$ moments, according to (8). In Eq. (10), the T- and lattice sum matrices read as
$\mathbf{T}(\omega)=\begin{pmatrix}T_{\rm MM}&T_{\rm MQ}\\ T_{\rm QM}&T_{\rm QQ}\end{pmatrix}\,,\quad\mathbf{\Sigma}^{\prime}(\omega)=\begin{pmatrix}\Sigma^{\prime}_{\rm MM}&\Sigma^{\prime}_{\rm MQ}\\ \Sigma^{\prime}_{\rm MQ}&\Sigma^{\prime}_{\rm QQ}\end{pmatrix}\,$, (24)
where $T_{\rm MM},\penalty 10000\ T_{\rm MQ},\penalty 10000\ T_{\rm QM},$ and $T_{\rm QQ}$ [resp. $\Sigma^{\prime}_{\rm MM}\equiv\Sigma^{\prime}_{\rm 0,0,0,0}(\omega,\bm{0}),\penalty 10000\ \Sigma^{\prime}_{\rm MQ}\equiv\Sigma^{\prime}_{\rm 0,0,2,0}(\omega,\bm{0}),$ and $\Sigma^{\prime}_{\rm QQ}\equiv\Sigma^{\prime}_{\rm 2,0,2,0}(\omega,\bm{0})$] are couplings between the corresponding multipoles induced by the particle [resp. lattice]. The definition of $\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\mathbf{k}_{\parallel})$ is given by Eq. (42). An isolated sphere possesses no particle-induced monopole-quadrupole coupling $T_{\rm MQ}=T_{\rm QM}=0$. Therefore, the matrix $\mathbf{T}^{-1}_{\rm eff}(\omega)$ from Eq. (9), when only the even multipoles are considered, becomes
@@
where, for a sphere, $T_{\rm MM}\equiv\mathfrak{a}_{0}$ and $T_{\rm QQ}\equiv\mathfrak{a}_{2}$ (see Appendix C for the definition of $\mathfrak{a}_{\ell}$). Thus, for a 2D lattice of spheres, the monopole-quadrupole coupling is fully induced by the lattice. It can be considered as a channel that enables the formation of the BIC in the symmetric metasurface with two subsystems, the sublattices of monopoles and quadrupoles, which interact through the radiation continuum.
For even zonal multipoles, $\mathcal{P}_{+}=\mathcal{P}_{-}$ according to Eq. (19). The formation of a BIC corresponds to $\mathcal{P}_{+}=\mathcal{P}_{-}=0$, which implies the following relationship between the multipole moments
$\quad a^{\rm eff}_{\rm M}-\sqrt{5}a^{\rm eff}_{\rm Q}=0\,.$ (26)
@@
The BIC formation in this system occurs when two radiating multipoles (in our case, zonal) with the same parity interfere, creating a nonradiant state. This mechanism is only accessible for an acoustic resonator on an infinite lattice when sound radiation must be suppressed only in the vertical directions. Since the radiation patterns of the multipoles of a single resonator do not overlap [see Fig. 1(b)], in this case, it is impossible to suppress the omnidirectional sound radiation through destructive interference of two multipoles. Moreover, the BIC formation in a lattice due to this mechanism is possible only with at least two zonal multipoles taken into account. In contrast, a single-multipole approximation is insufficient to obtain a nontrivial solution (because a single zonal multipole always radiates sound along the vertical axis). Finally, we need to emphasize that the considered metasurface can also support the BIC formed by the odd zonal multipoles, by the dipole and octupole (see Appendix I). The odd BIC emerges close to the frequency of the octupole resonance in the single particle, similar to the even BIC, which occurs close to the single-particle quadrupole resonance.
III.3 Realization in the infinite lattice

### v4

@@ II.3 Selection rules for multipole coupling in metasurfaces
Further consideration is facilitated by the symmetry analysis and associated selection rules that constrain which multipoles interact in structures with a given symmetry [53, 64, 48], or in other words, which off-diagonal elements of $\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ in Eq. (10) do not vanish.
@@
Let us first analyze the acoustic coupling of scalar multipoles induced by a two-dimensional (2D) Bravais lattice with rotational symmetry of order $\mathfrak{n}_{\rm L}$, considering rotations by $2\pi/\mathfrak{n}_{\rm L}$ around the $z$ axis. We also assume that each unit cell contains only one sphere, for which the single-particle T-matrix is diagonal $T_{\ell,m,\ell^{\prime},m^{\prime}}=\mathfrak{a}_{\ell}\delta_{\ell,\ell^{\prime}}\delta_{m,m^{\prime}}$, where $\delta_{\ell,\ell^{\prime}}$ is the Kronecker symbol (see Appendix C for the definition of $\mathfrak{a}_{\ell}$). Hence, the diagonal elements of $\mathbf{T}^{-1}_{\rm eff}$ become $\left(T^{-1}_{\ell,m,\ell,m}-\Sigma^{\prime}_{\ell,m,\ell,m}\right)$, while the off-diagonal elements are $\left(-\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}\right)$ with $\ell\neq\ell^{\prime}$ and $m\neq m^{\prime}$. Thus, the metasurface has inversion symmetry (considering $\mathbf{r}\to-\mathbf{r}$) and rotational symmetry of order $\mathfrak{n}_{\rm L}$. Lattice-induced coupling between two different multipoles $(\ell,m)$ and $(\ell^{\prime},m^{\prime})$ is allowed ($\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}\neq 0$) for an arbitrary $\mathbf{k}_{\parallel}$ if $\left(\ell+\ell^{\prime}\right)=(m-m^{\prime})\ \text{mod}\ 2$ [65]. Under normal incidence $\mathbf{k}_{\parallel}=\bm{0}$, the following conditions must also be satisfied (see Appendix E):
$\ell=\ell^{\prime}\ \text{mod}\ 2\,,\quad\quad m=m^{\prime}\ \text{mod}\ \mathfrak{n}_{\rm L}\,.$ (12)
@@
The first condition states that the multipoles must have the same parity, which follows from the inversion symmetry of the metasurface, while the second condition follows from its rotational symmetry.
However, even though $\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}=0$, the particle-induced coupling due to off-diagonal entries of the T-matrix $T_{\ell,m,\ell^{\prime},m^{\prime}}\neq 0$ can be present for nonspherical acoustic resonators. Shortly, the inversion symmetry of the particle requires $\ell=\ell^{\prime}\ \text{mod}\ 2$, while the rotational symmetry of order $\mathfrak{n}_{\rm P}$ implies $m=m^{\prime}\ \text{mod}\ \mathfrak{n}_{\rm P}$ [thus, the unit cell has the rotational symmetry of order $\mathfrak{n}=\min\left(\mathfrak{n}_{\rm L},\mathfrak{n}_{\rm P}\right)$]. Note that resonators can have other symmetries (e.g., reflection), which determine additional selection rules but are not considered here. The full list of selection rules for single acoustic resonators of different symmetries can be found in Ref. 53.
@@ II.4 Reflection and transmission coefficients
Effective multipole moments (11) allow us to calculate the reflection and transmission coefficients of a metasurface for an acoustic plane wave. To do this, we consider the pressure field in the far-field zone for $z<0$ and $z>0$. The incident wave is a pressure plane wave $p_{\rm inc}(\omega;\mathbf{r})=p_{0}\mathrm{e}^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}$ with $\mathbf{k}=\mathbf{k}_{\parallel}+\sqrt{k^{2}-k_{\parallel}^{2}}\hat{\mathbf{z}}$ and $|\mathbf{k}|=k$. Using Eqs. (3) and (5) with multipole coefficients (11), one can write the total pressure after the summation over the lattice sites as
@@
Note that the second term in the brackets in Eq. (13) corresponds to the scattered field in the presence of an external field $p_{0}\neq 0$ and the field of a lattice eigenmode in the absence $p_{0}=0$ (see Sec. III.1).
If the lattice constant is smaller than the wavelength in the background medium with $c_{\rm b}$, $L<\lambda_{\rm b}$, lattice field sums (14) can be written in the far-field region ($r\gg\lambda_{\rm b}$) as (see Appendix F)
$\Sigma_{\ell,m}(\omega,\mathbf{k}_{\parallel};\mathbf{r})=\frac{2\pi}{k^{2}A}(-\mathrm{i})^{\ell}Y_{\ell,m}(\theta_{\mathbf{k}_{\pm}},\varphi_{\mathbf{k}_{\pm}})\mathrm{e}^{\mathrm{i}\mathbf{k}_{\pm}\cdot\mathbf{r}}\,$, (15)
where the wave vectors are $\mathbf{k}_{\pm}=\mathbf{k}_{\parallel}\pm\sqrt{k^{2}-k_{\parallel}^{2}}\hat{\mathbf{z}}$ for $z\gtrless 0$ and $A$ is the area of a lattice unit cell ($A=L^{2}$ for a square unit cell). By inserting Eq. (15) into Eq. (13), we obtain the pressure field reflection and transmission coefficients for a metasurface as
@@
The quantities in Eqs. (16a)-(17) depend on $\omega$ and $\mathbf{k}_{\parallel}$ due to $a_{\ell,m}^{\rm eff}\equiv a_{\ell,m}^{\rm eff}(\omega,\mathbf{k}_{\parallel})$. In Eq. (17), $\mathbf{k}_{+}$ points into the direction of propagation of the incident and transmitted plane waves, while $\mathbf{k}_{-}$ is the wave vector of the reflected wave. Thus, a subwavelength biperiodic lattice can radiate only in two directions $\mathbf{k}_{-}$ and $\mathbf{k}_{+}$ with the amplitude and phase dictated by $\mathcal{P}_{-}$ and $\mathcal{P}_{+}$, respectively. The response of the lattice is determined by the effective multipole moments, i.e., the renormalized response of one unit cell (11). Moreover, the lattice coupling does not change the angular dependence of the multipoles of a unit cell, given by $Y_{\ell,m}(\theta_{\mathbf{k}_{\pm}},\varphi_{\mathbf{k}_{\pm}})$, but only changes the multipole moments from Eqs. (7) to (11). Thus, one can control the radiation pattern of the unit cell by tailoring the values of the multipole coefficients. Note also that, in the absence of the incident field $\mathbf{b}=\bm{0}$, expressions (17) describe the far-field amplitudes of the scattered waves radiated by a lattice eigenmode (see Sec. III.1).
Finally, let us consider the case of normal incidence $\mathbf{k}_{\parallel}=\bm{0}$. For the arguments $\cos\theta_{\pm}=\pm 1$ corresponding to $\theta=0$ (+z+z direction) and $\theta=\pi$ (z-z direction), the associated Legendre polynomials (31) take the following values [32]:
$P^{m}_{\ell}(\pm 1)$ $=\begin{cases}0,&m\neq 0\,,\\ (\pm 1)^{\ell},&m=0\,.\end{cases}$ (18)
From Eq. (17), we see that for $\mathbf{k}_{\parallel}=\bm{0}$, the amplitudes of the scattered (radiated) fields separate into two subseries corresponding to even and odd multipoles, in full agreement with the first selection rule from Eq. (12),
@@
where $\mathcal{P}_{\pm}\equiv\mathcal{P}_{\pm}(\omega,\bm{0})$ and $a_{\ell,0}^{\rm eff}\equiv a_{\ell,0}^{\rm eff}(\omega,\bm{0})$ for compactness. Thus, the response of a metasurface at normal incidence can be considered as a superposition of the responses of two noninteracting sublattices, each of which includes either only even multipoles or only odd multipoles, respectively.
III Accidental bound states in the continuum
@@ III.1 Eigenmode and BIC condition
According to Cramer‘s theorem, nontrivial solutions to homogeneous Eq. (20) exist if the inverse effective T-matrix is singular, i.e., $\det\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})=0$ [66]. The singularity of the inverse effective T-matrix indicates that the system has reached a resonance mode where a self-sustaining solution can exist without external stimulus.
For an open system and real $\mathbf{k}_{\parallel}$, a nontrivial solution to Eq. (20) can generally be found for a complex frequency $\omega(\mathbf{k}_{\parallel})-\mathrm{i}\gamma_{\rm r}(\mathbf{k}_{\parallel})-\mathrm{i}\gamma_{\rm nr}(\mathbf{k}_{\parallel})]$, where $\gamma_{\rm r}(\mathbf{k}_{\parallel})$ and $\gamma_{\rm nr}(\mathbf{k}_{\parallel})$ define energy dissipation due to radiative and nonradiative losses, respectively. However, photonic and acoustic metasurfaces can support eigenmodes without radiative losses $(\gamma_{\rm r}=0)$, which are just called bound states in the continuum or BICs. Moreover, in the absence of nonradiative losses (absorption), BICs emerge for pure real-valued frequencies due to $\gamma_{\rm nr}=0$. Hence, a criterion of the BIC occurrence can be formulated as [26, 67]
$\det\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})=0\,,\quad\quad\omega\in\mathbb{R}\,.$ (21)
Since the matrix $\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ can be subject to the singular value decomposition (SVD), the absolute value of its determinant is equal to the product of its singular values, which are nonnegative real-valued numbers [68]. Thus, if a singular value of the matrix is zero, then the determinant is also zero. Hence, the solution to Eq. (20) is the right singular vector of matrix $\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ with a singular value equal to zero [62, 69]. After denoting the operator that picks the smallest singular value of a matrix with $s_{\rm min}$, BIC condition (21) reads as
$s_{\rm min}\left[\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})\right]=0\,,\quad\quad\omega\in\mathbb{R}\,.$ (22)
Conditions (21) and (22) ensure the existence of a lattice eigenmode with a real eigenfrequency, i.e., a BIC. If this condition is satisfied, the right singular vector $\mathbf{a}^{\rm eff}(\omega,\mathbf{k}_{\parallel})$ corresponding to the eigenfrequency $\omega=\omega_{\rm BIC}$ contains the multipole coefficients $a_{\ell,m}^{\rm eff}(\omega,\mathbf{k}_{\parallel})$. Inserting them into Eq. (17) or (19) must give us that the far-field radiation amplitudes for the BIC are $\mathcal{P}_{+}=\mathcal{P}_{-}=0$. It is worth mentioning that criteria (21) and (22) are directly applicable to optical BICs with the corresponding change of the effective T-matrix.
@@ III.2 Even BIC in the monopole-quadrupole approximation
Equation (22) can provide us with resonance frequencies of BICs in arbitrary systems. Here, we give an example of the matrix $\mathbf{T}^{-1}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ for which Eq. (22) has a solution with $\mathbf{k}_{\parallel}=\bm{0}$ and a real eigenfrequency. For this mode, acoustic radiation in the $\pm z$ directions is suppressed, so that the metasurface does not radiate at all, and the mode can be referred to as an accidental BIC at the $\Gamma$-point. According to selection rules (12), eigenmodes (including BICs) of a lattice of spheres can have either even or odd parity, which also determines their multipolar content. In the following, we choose the case of even parity. However, the odd-parity modes, which can also enable an accidental BIC, can be considered in the same manner (see Appendix I).
Let us consider an eigenmode with even parity of the square lattice $\mathfrak{n}_{\rm L}=4$ in Eq. (12)] of spheres, for which the following multipole coefficients can be nonzero:
@@
where $\ell^{\prime}\in\mathds{N}_{0}$ and $m^{\prime}\in\mathds{Z}$, such that $|2m^{\prime}|\leq\ell^{\prime}$. The other multipole contributions must be zero according to selection rules (12). Furthermore, we also consider spherical scatterers (resonators) whose acoustic response in the given frequency range can be sufficiently described by the zonal monopole (M), dipole (D), quadrupole (Q), and octupole (O). At the same time, higher-degree multipoles can be neglected (see Fig. 2).
@@
In this case, we can keep only two terms from Eq. (23), the zonal monopole $a_{\rm M}^{\rm eff}\equiv a_{0,0}^{\rm eff}$ and the zonal quadrupole $a_{\rm Q}^{\rm eff}\equiv a_{2,0}^{\rm eff}$. For $\mathbf{k}_{\parallel}=\bm{0}$, all unit cells in the array will have identical zonal monopole $a_{\rm M}^{\rm eff}$ and zonal quadrupole $a_{\rm Q}^{\rm eff}$ moments, according to Eq. (8). In Eq. (10), the parts of the single-particle T-matrix and the lattice sum matrix corresponding to the zonal monopole and the zonal quadrupole read as
$\mathbf{T}(\omega)=\begin{pmatrix}T_{\rm MM}&T_{\rm MQ}\\ T_{\rm QM}&T_{\rm QQ}\end{pmatrix}\,,\quad\quad\mathbf{\Sigma}^{\prime}(\omega)=\begin{pmatrix}\Sigma^{\prime}_{\rm MM}&\Sigma^{\prime}_{\rm MQ}\\ \Sigma^{\prime}_{\rm MQ}&\Sigma^{\prime}_{\rm QQ}\end{pmatrix}\,$, (24)
where $T_{\rm MM},\penalty 10000\ T_{\rm MQ},\penalty 10000\ T_{\rm QM},$ and $T_{\rm QQ}$ [resp. $\Sigma^{\prime}_{\rm MM}\equiv\Sigma^{\prime}_{\rm 0,0,0,0}(\omega,\bm{0}),\penalty 10000\ \Sigma^{\prime}_{\rm MQ}\equiv\Sigma^{\prime}_{\rm 0,0,2,0}(\omega,\bm{0}),$ and $\Sigma^{\prime}_{\rm QQ}\equiv\Sigma^{\prime}_{\rm 2,0,2,0}(\omega,\bm{0})$] are couplings between the corresponding multipoles induced by the particle [resp. lattice]. The definition of $\Sigma^{\prime}_{\ell,m,\ell^{\prime},m^{\prime}}(\omega,\mathbf{k}_{\parallel})$ is given by Eq. (43). An isolated sphere does not possess particle-induced coupling; therefore, $T_{\rm MQ}=T_{\rm QM}=0$. Therefore, the matrix $\mathbf{T}^{-1}_{\rm eff}(\omega)$ from Eq. (9), when only the even multipoles are considered, becomes
@@
where, for a sphere, $T_{\rm MM}\equiv\mathfrak{a}_{0}$ and $T_{\rm QQ}\equiv\mathfrak{a}_{2}$ (see Appendix C for the definition of $\mathfrak{a}_{\ell}$). Thus, for a biperiodic lattice of spheres, the monopole-quadrupole coupling is fully induced by the lattice. It can be considered as a channel that enables the formation of the BIC in the symmetric metasurface with two subsystems, the sublattices of monopoles and quadrupoles.
For even zonal multipoles, $\mathcal{P}_{+}=\mathcal{P}_{-}$ according to Eq. (19). The formation of a BIC corresponds to $\mathcal{P}_{+}=\mathcal{P}_{-}=0$, which implies the following relationship between the multipole moments of the even BIC:
$a^{\rm eff}_{\rm M}-\sqrt{5}a^{\rm eff}_{\rm Q}=0\,.$ (26)
@@
The BIC formation in this system occurs when two radiating multipoles (in our case, zonal) with the same parity interfere, creating a nonradiant state. This mechanism is accessible for an acoustic resonator on an infinite lattice only when acoustic radiation must be suppressed solely in the normal directions. Since the radiation patterns of the multipoles do not overlap [see Fig. 1(b)], it is impossible in this case to suppress the omnidirectional acoustic radiation for a single resonator through destructive interference between two multipoles. Moreover, BIC formation in a metasurface via this mechanism is possible only when at least two zonal multipoles are taken into account. In contrast, a single-multipole approximation is insufficient to obtain a nontrivial solution, because a single zonal multipole always radiates acoustic waves along its axis.
It is worth mentioning that the considered metasurface can also support an odd BIC formed due to destructive interference between the dipole and the octupole that satisfy the following constraint (see Appendix I)
$a^{\rm eff}_{\rm D}-\sqrt{7/3}a^{\rm eff}_{\rm O}=0.$ (27)
The odd BIC emerges close to the frequency of the octupole resonance of a single scatterer, similar to the even BIC, which occurs close to the quadrupole resonance of a single scatterer.
Finally, let us discuss the impact of lattice and resonator symmetries. The form of matrix (25) remains the same for the hexagonal lattice $\mathfrak{n}_{\rm L}=6$ in Eq. (12)], but with different values of the lattice sums. The monoclinic and orthorhombic lattices ($\mathfrak{n}_{\rm L}=2$) enable the coupling between the zonal multipoles and the multipoles with $m=\pm 2$, which, however, do not radiate acoustic waves in the normal direction [see Eq. (19)]. For a nonspherical resonator on a metasurface, the eigenmodes, including BICs, can be classified by parity as long as the metasurface possesses inversion symmetry. In its absence, the zonal even and odd multipoles are coupled, and the genuine BICs with even or odd parity transform into quasi-BICs for the parameters when Eq. (26) or (27) is satisfied, respectively (small shape asymmetries of the resonators, which can occur in reality, will also lead to a decrease in the radiative $Q$ factor). To establish a genuine BIC in a metasurface without inversion symmetry, the multipole moments must simultaneously satisfy Eqs. (26) and (27). In contrast, breaking rotational symmetry of the resonator with respect to the $z$ axis while preserving inversion symmetry induces only coupling between zonal and nonzonal multipoles with the same parity [53], which does not affect Eqs. (26) and (27); therefore, genuine BICs with a certain parity can exist in such metasurfaces. Note that additional off-diagonal terms due to $T_{\rm MQ}=T_{\rm QM}\neq 0$ appear in Eq. (25) for nonspherical particles.
III.3 Realization in an infinite metasurface